) High risk funds of the $50,000 that Natasha pocketed on her last real estate deal, 20,000 went to charity. She invested part of the remainder in Dreyfus New Leaders Fund with an annual yield of 16% and the rest in Templton Growth fund with an annual yield of 25% if she made $6060 on these investments in one year, then how much did she invest in each fund?
Answer by stanbon(75887)

High risk funds of the $50,000 that Natasha pocketed on her last real estate deal, 20,000 went to charity. She invested part of the remainder in Dreyfus New Leaders Fund with an annual yield of 16% and the rest in Templton Growth fund with an annual yield of 25% if she made $6060 on these investments in one year, then how much did she invest in each fund?
---------------------
She invested 50k - 20k = $30,000
--------------
Let amount invested at 16% be "x"; interest on this is 0.16x dollars
-----------------
Amount invested at 25% is "30,000-x" ; interest on this is 0.25(30,000-x)
=7500-0.25x dollars.
------------------------
EQUATION:
interest + interest = 6060 dollars
0.16x + 7500-0.25x = 6060
-0.09x = -1440
x=$16000 (amount invested at 16%)
30000-x = $14000 (amount invested at 25%)
================
Cheers,
Stan H.